### v3

# UniVST: A Unified Framework for Training-free Localized Video Style Transfer

Abstract
This paper presents UniVST, a unified framework for localized video style transfer based on diffusion model. It operates without the need for training, offering a distinct advantage over existing diffusion methods that transfer style across entire videos. The endeavors of this paper comprise: (1) A point-matching mask propagation strategy that leverages the feature maps from the DDIM inversion. This streamlines the model’s architecture by obviating the need for tracking models. (2) A training-free AdaIN-guided video style transfer mechanism that operates at both the latent and attention levels. This balances content fidelity and style richness, mitigating the loss of localized details commonly associated with direct video stylization. (3) A sliding-window consistent smoothing scheme that harnesses optical flow within the pixel representation and refines predicted noise to update the latent space. This significantly enhances temporal consistency and diminishes artifacts in stylized video. Our proposed UniVST has been validated to be superior to existing methods in quantitative and qualitative metrics. It adeptly addresses the challenges of preserving the primary object’s style while ensuring temporal consistency and detail preservation. Our code is available at https://github.com/QuanjianSong/UniVST.
Index Terms:
Diffusion Model, Training Free, Video Style Transfer.
1 Introduction
Video editing has greatly improved thanks to the use of diffusion methods [1, 2, 3]. T2V-Zero [4] changes self-attention mechanisms to cross-frame attention. Tune-A-Video customizes video editing by adjusting attention weights [5]. Fate-Zero [6] keeps the video content intact via information from the inversion process. Animate-Zero [7] and VideoBooth [8] add layers that stress time.
Figure 1: Existing methods suffer from (a) lack of fine-grained control, (b) balance between content fidelity and style richness, and (c) temporal inconsistency. The last row shows the visualization of temporal consistency using MimicMotion’s Y-T slices [9] for both the existing method [10] and our proposed UniVST.
Video stylization, a sub-area of video editing, is gaining popularity. It is about adding artistic styles to video content. Our review identifies two main video stylization methods. The first method uses image or text to direct the style. StyleCrafter [11] personalizes video style with extra training adapter; VideoBooth [8] uses image for custom video creation prompts; while Diffutoon [12] focuses on cartoon coloring from text descriptions. The second method includes all-encompassing video editing frameworks like AnyV2V [13] and BIVDiff [10] that can perform various tasks, including style transfer. These usually apply a consistent style across the whole video, good for thematic expression but sometimes lacking in precision for specific uses. For example, as shown in Fig. 1, in film, scenes might be shot normally and later stylized to match different environments. This enables aried scene changes from one video take, which is also useful in advertising for efficiently applying local style changes to existing videos. Therefore, selectively styling parts of a video while keeping others unchanged is an important academic and commercial task.
We illustrate Fig. 1 and pinpoint three major limitations of current methods: 1. Lack of Fine-grained Control. While some existing approaches [6, 5] leverage localized text descriptions to preserve the overall style transfer and maintain the primary object’s style, the control is usually too coarse. This lack of precision can lead to the model failing to understand the nuances, potentially causing unintended style transfer effects on objects not explicitly mentioned in the text. 2. Content Fidelity vs. Style Richness. In the realm of video stylization, there is a delicate balance to be struck between content fidelity and style richness. An overemphasis on style richness can result in a blurring and alter the original layout of the video, whereas an excessive focus on content fidelity might yield a stylized outcome that is indistinguishable from the original. 3. Temporal Inconsistency. Unlike image stylization, video stylization requires careful consideration of temporal coherence between frames. As seen in [10, 13], merely extending image stylization techniques to video can lead to inconsistencies across frames, manifesting as flickering and artifacts.
In this paper, we propose UniVST, solving the above issues in a unified framework for localized [^1] Here, “localized” indicates a particular part of the video. video style transfer in a training-free manner. We inject the latent information from the content branch into the editing branch using mask, enabling the model to preserve the primary object’s style. Unlike in the image domain, where masks can be readily obtained using additional models such as those in [14, 15], the video domain presents a more labor-intensive challenge due to the need for frame-by-frame mask generation. To address this, we introduce a point-matching mask propagation strategy leveraging feature maps from the DDIM inversion process [16] to capture correlations. Inspired by the image stylization model [17], replacing the key and value in the editing branch with those from style branch can facilitate style transfer. However, we found that directly applying this method in video style transfer can lead to the loss and blurring of local details. Based on this foundation, we develop a training-free AdaIN-guided style transfer mechanism, which functions at both the latent and attention levels. It adeptly balances content fidelity with style richness. Detailed comparison of results is presented in Fig. 4.
Our survey shows that existing methods often utilize optical flow information to achieve temporal consistency. For instance, Ground-A-Video [2] uses optical flow to smooth initial noise, ensuring its temporal coherence, while Flatten [1] integrates optical flow into the attention process to maintain frame consistency. Building on these insights, we present a sliding-window consistent smoothing scheme that employs optical flow within the pixel representation and refines predicted noise to update the latent space. As illustrated in Fig. 1, our method reduce flickering and artifacts while enhances the temporal consistency of the edited video. Overall, our major contributions in this paper are as follows:
  • We introduce, to the best of our knowledge, the first localized video style transfer framework, featuring a novel point-matching mask propagation strategy.
  • We develop a training-free AdaIN-guided video style transfer mechanism that operates at both the latent and attention levels. It effectively harmonizes content fidelity and style richness throughout the transfer process.
  • We present a sliding-window consistent smoothing scheme based on optical flow, which adeptly upholds temporal consistency during the video transfer process.
  • Extensive experiments demonstrate that our framework outperforms several state-of-the-art methods in both qualitative and quantitative metrics.
2 Related Work
Image Style Transfer. RSCT [18] calls for the separate foreground and background style transfer, and combines them to ensure a consistent style. DiffStyle [19] employs LoRA [20] to refine image prompts for stylization, while ArtBank [21] enhances text embeddings through fine-tuning image prompts. StyleID [17] leverages key-value replacement coupled with an initial latent AdaIN [22] to preserve content fidelity. Z* [23] rearranges attention to integrate content with style. InstanceStylePlus [24] utilizes adapters to safeguard content retention and enhance stylistic expression.
Video Style Transfer. OCD [25] integrates object-centric sampling and merging, expediting the editing process. StyleCrafter [11] refines adapters using image prompts based on the T2V model. Style-A-Video [26] enhances the stylistic alignment through strategic style guidance. Diffutoon [12] pioneers primary and editing branches for multi-segment editing in video colorization, employing FastBlend [27] to uphold temporal continuity. BIVDiff [10] combines frame-by-frame image editing with temporal consistency modeling. AnyV2V [13] follows a similar initial editing strategy, subsequently leveraging an I2V model [28] to propagate changes across the video sequence.
Figure 2: Overall framework. It is structured around three main components: (1) Point-Matching Mask Propagation, (2) AdaIN-Guided Video Style Transfer(Attention-shift and latent-shift) and (3) Sliding-Window Consistent Smoothing.
3 Preliminary
Latent Diffusion Model. LDM [29] adds and removes noise in a low-dim space by using the encoder $\mathcal{E}$ and decoder $\mathcal{D}$. Given a Gaussian noise $Z_{T}$, DDIM denoising can be formulated as [3]:
$Z_{t-1}=\sqrt{\alpha_{t-1}}\underbrace{{Z}_{t\rightarrow 0}}_{\text{predicted }Z_{0}}+\underbrace{\sqrt{1-\alpha_{t-1}}\epsilon_{\theta}\left(Z_{t},t,C\right)}_{\text{direction pointing to }Z_{t-1}},$ (1)
where ${Z}_{t\rightarrow 0}$ is an estimation of $Z_{0}$ at time step $t$:
${Z}_{t\rightarrow 0}=\left(Z_{t}-\sqrt{1-{\alpha}_{t}}\epsilon_{\theta}\left(Z_{t},t,C\right)\right)/\sqrt{\alpha_{t}},$ (2)
where $\alpha_{t}$ is a parameter and $C=\phi$ for editing integrity.
DDIM Inversion. DDIM [16] transforms noise $Z_{T}$ into $Z_{0}$. Given reversible ODE [30], its reverse process can be described by:
${Z}_{t+1}={A}_{t}{Z}_{t}+{B}_{t}\epsilon_{\theta}\left({Z}_{t},t,\phi\right),$ (3)
where $A_{t}$ and $B_{t}$ are functions of time $t$. During editing, the initial noise $Z_{T}$ that aligns with the original video distribution $Z_{0}$ can be obtained using this equation.
3D UNet Extension. For video tasks, LDM requires a 3D UNet extension. Following [31, 1], the $3\times 3$ convolutional kernel in the convolutional blocks is expanded to $1\times 3\times 3$. To achieve frame parallelization, the feature shape $Z_{t}$ is transformed from $\mathbb{R}^{b\times c\times f\times h\times w}$ to $\mathbb{R}^{(bf)\times c\times h\times w}$ before self-attention and cross-attention operations. To facilitate frame interaction, we follow T2V-Zero [4] to modify the self-attention to cross-frame attention :
$\begin{array}[]{c}Q^{i}=W^{query}Z^{i},\quad K^{i}=W^{key}\left[Z^{{1}},Z^{{i-1}}\right],\\ V^{i}=W^{value}\left[Z^{{1}},Z^{{i-1}}\right],\end{array}$ (4)
where $Z^{i}$ denotes the feature of the $i$-th frame, and $W^{query}$, $W^{key}$, $W^{value}$ represent the respective mapping matrices, with [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] indicating concatenation.
Algorithm 1 Point-Matching Mask Propagation
1:Input: $M_{1}$ Mask for the first frame.
2:Input: $F_{t_{0}}$ Feature map from upsampling block-2.
3:Input: $r$ The sampling rates.
4:Input: $k$ The $k$-nearest points.
5:Input: $n$ The number of reference frames.
6:Output: $\{M_{i}\}_{i=1}^{N}$ Masks for all frames.
7:Initialize: previous_features [$\leftarrow$[\,]← [ ] Initialize empty list for previous features.
8:Initialize: previous_masks [$\leftarrow$[\,]← [ ] Initialize empty list for previous masks.
9:Initialize: first_feature [$\leftarrow$[\,]← [ ] Initialize empty list for first feature.
10:for $i=1,\dots,N$ do
11:    $fore\_index\leftarrow\text{where}(M_{i-1}=1)$
12:    $back\_index\leftarrow\text{where}(M_{i-1}=0)$
13:    $fore\_index\leftarrow\text{random\_sample}(fore\_index,r\cdot\frac{\lvert\text{fore}\rvert}{\lvert\text{fore}\rvert+\lvert\text{back}\rvert})$
14:    $back\_index\leftarrow\text{random\_sample}(back\_index,r\cdot\frac{\lvert\text{back}\rvert}{\lvert\text{fore}\rvert+\lvert\text{back}\rvert})$
15:    $current\_index\leftarrow\text{concat}(fore\_index,back\_index)$
16:    $current\_feature\leftarrow F_{t_{0}}^{i}[current\_index]$
17:    if $i=1$ then
18:       first_feature.append($current\_feature$) Store first frame.
19:       $M_{i}\leftarrow M_{1}$
20:    else
[…]
22:          previous_features.pop(0) Remove the oldest frame feature.
23:          previous_masks.pop(0) Remove the oldest frame mask.
24:       end if
[…]
26:       $M_{i}\leftarrow\text{KNN}(f,F_{t_{0}}^{i},\text{concat}(\text{previous\_masks},M_{1}),k)$
27:       previous_features.append($current\_feature$) Store current frame.
28:       previous_masks.append($M_{i}$) Store current mask.
29:    end if
30:end for
31:Return: $\{M_{i}\}_{i=1}^{N}$
4 Localized Video Style Transfer
Given an original video comprising $N$ frames $\{I^{i}\}_{i=1}^{N}$ and a reference style image $I^{s}$, our objective is to transform the original video into a new sequence $\{\mathcal{I}^{i}\}_{i=1}^{N}$. To achieve this, we propose UniVST, which integrates the style of the reference image into the original video while preserving the style of the primary objects within the video unchanged. As shown in Fig. 2, it mainly includes Point-Matching Mask Propagation, AdaIN-Guided Video Style Transfer, and Sliding-Window Consistent Smoothing three components.
4.1 Point-Matching Mask Propagation
To preserve the primary style during the style transfer process, a potent technique is to apply a mask to blend the latent variables at several steps. This can be expressed as:
$\mathcal{Z}_{t}=M\cdot Z_{t}+(1-M)\cdot\mathcal{Z}_{t},$ (5)
where $\mathcal{Z}_{t}$ and $Z_{t}$ denote the latent variables in the editing and content branch, respectively. However, in contrast to images, generating masks $M$ for each video frame is laborious. Users typically prefer to control the video’s main style by supplying a mask for only the initial frame. Extending the user-provided mask across all frames efficiently is necessary. Thus, we introduce point-matching mask propagation below and details are provided in Algorithm 1.
Figure 3: Comparison of accuracy and inference time under different mask propagation strategies. The introduction of anchor frames with the downsampling strategy significantly mitigates the accuracy degradation in subsequent frame propagation and also reduces inference time.
Point Matching. Inspired from DIFT [32], the three upsampling blocks in the UNet are replete with rich semantic information. Therefore, we can retain the feature map $\{F_{t_{0}}^{i}\}_{i=1}^{N}$ from upsampling block-2 at a given step $t_{0}$ during DDIM inversion in the source video, and use them to calculate the pixel correspondences between frames. Specifically, to pinpoint the pixel in the $j$-th frame that most closely resembles a given point $p_{i}$ in the $i$-th frame, we calculate the cosine similarity between the corresponding point in the feature map $F_{t_{0}}^{i}$ and all points in the feature map $F_{t_{0}}^{j}$ as:
$p_{j}={\underset{p_{j}\in F_{t_{0}}^{j}}{\arg\min}}\ CosSim\big{(}p_{i},p_{j}\big{)},\quad p_{i}\in F_{t_{0}}^{i}.$ (6)
Utilizing this principle, we can propagate the mask from the first frame to all subsequent frames using the mask propagation strategy formally introduced in the following.
Mask Propagation. Given an initial object mask for the first frame from the user, we use $k$-NN [33] to propagate this mask to subsequent frames. For each point in the $i$-th frame, we find the $k$ most similar points in the $(i-1)$-th frame using point matching. If the majority of these $k$ points are in the foreground, the corresponding point in the $i$-th frame is also classified as foreground; otherwise, it is background.
Figure 4: (a) Original video frame; (b) Key-value replacement [17]; (c) Key-value AdaIN; (d) Key-value AdaIN with gradual attention shift.
However, as shown in Fig. 3, using a naive autoregressive approach for mask propagation can lead to error accumulation, significantly decreasing accuracy for subsequent frames. To address this, we design the anchor frames mechanism that incorporates information from the first frame and the previous $n$ frames.
Although introducing such a mechanism can improve the propagation accuracy of subsequent frames, it requires more similarity calculations, potentially reducing efficiency. To address this issue, we apply random downsampling at a rate of $r$ to both foreground and background regions of the anchor frames to minimize costs. The sampling rates are then adjusted to reflect the proportions in the foreground and background. As illustrated in Fig. 3, the combination of anchor frames and random downsampling maintains final mask propagation accuracy while improving efficiency. More comprehensive information and in-depth analysis can be available in the experimental section.
4.2 AdaIN-Guided Video Style Transfer
By using DDIM inversion [34] to the original video $\{I^{i}\}_{i=1}^{N}$ and the reference style image $I^{s}$, we obtain their respective noise latents $Z_{t}$ and $Z_{t}^{s}$ ($t=1\rightarrow T$). Then, we establish the edited latents $\mathcal{Z}_{t}=Z_{t}$, and integrate a three-branch architecture with latent-shift and attention-shift mechanisms to accomplish video style transfer.
AdaIN-Guided Latent Shift. We have discovered that applying the AdaIN [17] to the initial noise significantly enhances the transferred color style. Therefore, we extend the application of AdaIN to several steps during the denoising process. Within the interval $t\in[\tau_{0},\tau_{1}]$, We consider applying it to the latent $\mathcal{Z}_{t}$ of the editing branch, shifting its mean and variance to the latent $Z_{t}^{s}$ of the style branch:
$\mathcal{Z}_{t}$ $=\text{AdaIN}(\mathcal{Z}_{t},Z_{t}^{s}),$ (7)
$\text{AdaIN}(\mathcal{Z}_{t},Z_{t}^{s})=\sigma\left(Z_{t}^{s}\right)\left(\frac{\mathcal{Z}_{t}-\mu\left(\mathcal{Z}_{t}\right)}{\sigma\left(\mathcal{Z}_{t}\right)}\right)+\mu\left(Z_{t}^{s}\right),$ (8)
where $\mu(\cdot)$ and $\sigma(\cdot)$ denoting mean and standard deviation.
AdaIN-Guided Attention Shift. To prevent the direct replacement of self-attention from causing content damage, [17] blended the original video’s query $Q_{t}$ with the edited query $\mathcal{Q}_{t}$ as:
$\begin{split}\tilde{\mathcal{Q}}_{t}&=\gamma\cdot\mathcal{Q}_{t}+(1-\gamma)\cdot Q_{t},\\ &\tilde{\mathcal{K}}_{t}=K_{t}^{s},\;\tilde{\mathcal{V}}_{t}=V_{t}^{s},\end{split}$ (9)
where $\gamma$ controls the balance between content and style. However, as observed in Fig. 4(b), this approach still lead to a loss of localized details in video stylization. To mitigate the loss of details caused by direct replacement, we propose to shift the mean and variance of the distributions of $\mathcal{K}_{t}$ and $\mathcal{V}_{t}$ using AdaIN:
$\tilde{\mathcal{K}}_{t}=\text{AdaIN}(\mathcal{K}_{t},K_{t}^{s}),\;\tilde{\mathcal{V}}_{t}=\text{AdaIN}(\mathcal{V}_{t},V_{t}^{s}).$ (10)
Though this shift preserves localized details, as indicated in Fig. 4(c), it causes content blurring and artifacts. Upon analysis, we find that the early denoising stages contain less information, making key and value replacements inappropriate. Thus, within the interval $t\in[\tau_{2},\tau_{3}]$, we gradually decrease the impact of the attention shift by combining the key and value replacements with the query blending:
$\tilde{\mathcal{K}}_{t}$ $=\beta_{t}\cdot\text{AdaIN}(\mathcal{K}_{t},K_{t}^{s})+(1-\beta_{t})\cdot K_{t}^{s},$ (11)
$\tilde{\mathcal{V}}_{t}$ $=\beta_{t}\cdot\text{AdaIN}(\mathcal{V}_{t},V_{t}^{s})+(1-\beta_{t})\cdot V_{t}^{s},$
$\beta_{t}$ $=\frac{\beta_{\tau_{3}}-\beta_{\tau_{2}}}{\tau_{3}-\tau_{2}}\cdot(t-\tau_{2})+\beta_{\tau_{2}}.$
By combing key-value AdaIN with gradual attention shift in Eq. (11) and query blending in Eq. (9), the stylization results achieve an effective balance between content fidelity and style richness, as shown in Fig. 4(d) and Table VII.
4.3 Sliding-Window Consistent Smoothing
We address the common issues of flicker and artifacts in video stylization by focusing on temporal inconsistency. Following [1, 2], we utilize optical flow with specialized sliding window smoothing.
Optical Flow Warping. Optical flow estimation is a pivotal technique, often used for object tracking [35, 36]. It predicts the motion direction and velocity of pixels by analyzing changes between consecutive frames. Using models like RAFT [37], optical flow information can be predicted from one frame to the next, allowing for the reconstruction of the initial frame. In cases where occlusion occurs, bidirectional optical flow can estimate the mask and fill in occluded regions with the original image. This process, termed “Warp,” is key to our following discussions.
Sliding Window Smoothing. We apply the warp operation during denoising process to achieve local smoothing. At time step $t$, we calculate the latent representation $\mathcal{Z}_{t\rightarrow 0}$ from the predicted noise $\varepsilon_{t}$ according to Eq. (2). $\mathcal{Z}_{t\rightarrow 0}$ is then decoded by the decoder $\mathcal{D}$ into a pixel representation $\{P_{t}^{i}\}_{i=1}^{N}$. For each decoded frame $P_{t}^{i}$, we apply a sliding window of size $2m$ to warp its neighboring frames and smooth them to update the original pixel:
$\bar{P}_{t}^{i}\leftarrow\frac{1}{2m+1}\cdot\sum_{j=i-m}^{i+m}\operatorname{Warp}(P_{t}^{i},P_{t}^{j}).$ (12)
The updated values from the current window will be incorporated into the calculations for the next window, as shown in Fig. 2. Subsequently, the VAE’s encoder $\mathcal{E}$ re-encodes the smoothed $\{\bar{P}_{t}^{i}\}_{i=1}^{N}$ into latent representations $\bar{\mathcal{Z}}_{t\rightarrow 0}$, which is an updated version of $\mathcal{Z}_{t\rightarrow 0}$. Next, the predicted noise $\varepsilon_{t}$ is adjusted using the inversion of Eq. (2):
$\bar{\varepsilon}_{t}\leftarrow\frac{\mathcal{Z}_{t}-\sqrt{\alpha_{t}}\bar{\mathcal{Z}}_{t\rightarrow 0}}{\sqrt{1-\alpha_{t}}}.$ (13)
Finally, the latent representation is updated in accordance with the DDIM schedule in Eq. (1):
$\mathcal{Z}_{t-1}\leftarrow\sqrt{\alpha_{t-1}}\bar{\mathcal{Z}}_{t\rightarrow 0}+\sqrt{1-\alpha_{t-1}}\bar{\varepsilon}_{t}.$ (14)
It is important to note that we do not adopt a global smoothing strategy because extensive optical flow predictions would introduce significant computational overhead. Instead, by employing a sliding local window scheme, we achieved a balance between temporal consistency and efficiency. And this scheme is only performed within the interval $t\in[\tau_{4},\tau_{5}]$.
TABLE I: Quantitative comparison of existing methods across various metrics, superscript * indicates non-diffusion methods. The bold values represent the best performance and the underscore stresses the second best.
Datasets Methods Overall Transfer Performance Style Preservation Consistency Efficiency
[…]
DAVTG-WikiArt Diffutoon [12] 43.150 26.268 0.582 0.961 0.954 0.980 44.263
StyleCrafter [11] 39.119 21.008 0.777 0.928 0.895 0.978 111.225
AnyV2V [13] 33.774 19.104 0.680 0.938 0.914 0.978 78.474
BIVDiff [10] 46.071 25.468 0.740 0.921 0.894 0.955 158.367
EFDM [38] 37.407 22.350 0.602 0.936 0.935 0.963 5.969
CAST [39] 45.198 26.174 0.663 0.948 0.936 0.975 7.348
UniVST (Ours) 37.152 26.134 0.369 0.986 0.991 0.980 162.745
DAVTG-Laion Diffutoon [12] 38.610 23.700 0.563 0.959 0.953 0.979 44.263
StyleCrafter [11] 43.161 23.246 0.780 0.926 0.897 0.978 111.225
AnyV2V [13] 35.648 20.019 0.696 0.936 0.918 0.975 78.474
BIVDiff [10] 36.676 20.156 0.733 0.921 0.897 0.956 158.367
EFDM [38] 35.673 21.549 0.582 0.933 0.931 0.964 5.969
CAST [39] 34.828 19.974 0.661 0.947 0.936 0.968 7.348
UniVST (Ours) 30.636 21.146 0.383 0.986 0.990 0.981 162.745
5 Experimentation
5.1 Experimental Settings
Our method is based on the Stable Diffusion V1.5 model, as referenced in the literature [29]. The hyperparameters are configured as follows: $\tau_{0}=0.1T$, $\tau_{1}=0.2T$, $\tau_{2}=0.4T$, $\tau_{3}=1.0T$, $\tau_{4}=0.5T$, $\tau_{5}=0.6T$, $t_{0}=0.4T$, with the total time period $T$ set to 50. Additionally, we have $\gamma=0.35$, $\beta_{\tau_{2}}=0.1$, $\beta_{\tau_{3}}=0.9$, $r=0.3$, $k=15$, $m=2$, and $n=9$. Videos are resized to a resolution of 512 by 512 pixels and are processed in batches consisting of 16 frames each. For optical flow estimation, we employ the RAFT model [37]. Our approach does not require training or fine-tuning and can be executed on a single RTX 3090 GPU, making it both efficient and practical for deployment.
Datasets. Building upon the foundational work referenced in [17, 26], we have meticulously selected style images from two renowned databases: WikiArt [40] and Laion-Aesthetics-6.5+ [41], each contributing a collection of 20 distinct images. For the content videos, we have procured a dataset comprising 50 videos from the DAVIS2016 [42] and TGVE [43] datasets, which we have amalgamated and designated as DAVTG. Through the integration of these datasets, we have successfully established two comprehensive subsets: DAVTG-WikiArt, which contains 1,000 curated pairs, and DAVTG-Laion, which also features 1,000 paired entries.
Metrics. (1) Overall Transfer Performance: Drawing inspiration from the methodologies outlined in [17, 26], we assess the overall transfer performance by employing ArtFID [44], a metric that amalgamates LPIPS [45] for gauging content fidelity and FID [46] for evaluating the richness of style. (2) Foreground Style Preservation: To evaluate the preservation of texture, we utilize the SSIM [47] metric, while the CLIP-I score [48] is employed to assess the preservation of semantic content. During this evaluation, the background is neutralized to ensure accuracy. (3) Temporal Consistency: In line with the approaches detailed in [26, 11], we measure the similarity between consecutive frames using the CLIP-F score [48], which serves as a metric for evaluating the temporal coherence of the video. (4) Inference Efficiency: To gauge the time efficiency, we calculate the total inference time, thereby assessing their performance in terms of computational speed.
Figure 5: Qualitative comparison of our proposed UniVST with existing methods. superscript * indicates non-diffusion methods.
5.2 Quantitative Comparison
We have chosen a selection of video stylization techniques, including Diffutoon [12] and StyleCrafter [11], as well as two video editing frameworks: AnyV2V [13] and BIVDiff [10]. Additionally, we have considered two traditional non-diffusion methods: EFDM [38] and CAST [39]. Given that our specific task does not perfectly correspond with some of the existing methods, we have made necessary modifications to ensure a fair comparison. In the case of Diffutoon [12], which is designed to process text inputs for coloring tasks and does not natively support style images, we have overcome this limitation by leveraging Chat-GPT4 [49] to generate descriptive captions for the style images. For the style generation task with StyleCrafter [11], which does not inherently support video editing, we employ inversion to obtain the initial noise of the original video, which is then utilized in the generation process. For the unified video editing frameworks, AnyV2V [13] and BIVDiff [10], we have integrated InstantStyle [50] as the image editing model to enhance their capabilities. As for the traditional non-diffusion methods, EFDM [38] and CAST [39], which are well-aligned with our task, we have retained them in their original form without making any alterations.
We employ these six established methods as benchmarks to evaluate our proposed UniVST, with the quantitative comparisons presented in Table I. The results demonstrate that our UniVST outperforms the others across all three performance metrics. On both the DAVTG-WikiArt and DAVTG-Laion datasets, UniVST consistently surpasses the competition in terms of foreground style preservation and temporal consistency, showcasing its robust capability for localized video style transfer. Regarding overall transfer performance, UniVST claims the top position on the DAVTG-Laion dataset and secures the second spot on DAVTG-WikiArt. Although localized transfer tasks might marginally affect overall performance, leading to a slightly higher FID score compared to other methods, these findings underscore UniVST’s proficiency in effectively reconciling content fidelity with style richness.
Our inference efficiency does not stand out. This is because we opt not to adopt the temporal attention mechanisms in existing T2V frameworks [11, 26]. Instead, we utilize the Sliding-Window Consistent Smoothing strategy, which requires dynamic optical flow estimation during the denoising process. While this approach improves temporal consistency, it also leads to increased computational overhead. In Fig 13, enlarging the sliding window size results in a proportional increase in the number of frames needed for optical flow estimation, thereby raising the time cost. This trade-off between temporal consistency and computational efficiency highlights the potential optimization in future work.
Figure 6: Qualitative comparisons of our proposed UniVST with some existing commercial models. The first row presents the original video, reference image, and mask, followed by the results of Gen-1 [51] in the second row, Gen-3 in the third row, and our proposed UniVST in the final row.
TABLE II: Ablation study for proposed components. The bold values represent the best performance.
Datasets Methods Overall Transfer Performance Style Preservation Consistency
[…]
DAVTG-WikiArt UniVST 37.152 26.134 0.369 0.986 0.991 0.980
w/o mask-guidance 38.014 26.050 0.405 0.966 0.970 0.979
w/o latent-shift 37.649 26.909 0.349 0.985 0.990 0.980
w/o attention-shift 39.589 27.506 0.388 0.985 0.988 0.983
w/o window-smoothing 38.626 26.907 0.384 0.982 0.989 0.970
DAVTG-Laion UniVST 30.636 21.146 0.383 0.986 0.990 0.981
w/o mask-guidance 30.848 20.974 0.404 0.961 0.967 0.979
w/o latent-shift 31.001 21.766 0.362 0.985 0.989 0.980
w/o attention-shift 32.534 22.623 0.377 0.985 0.989 0.981
w/o window-smoothing 30.834 21.249 0.386 0.981 0.988 0.971
5.3 Qualitative Comparison
For a qualitative assessment, we have juxtaposed the editing results with those of six other baseline methods, as illustrated in Fig. 5. Diffutoon [12], which is renowned for its expertise in cartoon coloring, exhibits a limited responsiveness to a variety of image styles. This results in a final edited video that is notably monotonous, heavily inclined towards a cartoonish anime aesthetic. StyleCrafter [11], tailored for style-based video generation, demonstrates a high sensitivity to stylistic cues. However, it encounters difficulties in preserving content fidelity, leading to substantial distortion of background details and character features during the transfer process. BIVDiff [10] falls short in both content fidelity and temporal consistency, similarly altering background information and character traits during the transfer. AnyV2V [13] excels in maintaining temporal consistency but faces challenges with transfer tasks involving a central object. It often over-transfers the subject’s style, resulting in a loss of its distinctive attributes. The traditional non-diffusion methods, EFDM [38] and CAST [39], compromise detail information during the transfer process and introduce color artifacts. On the contrary, UniVST demonstrates a remarkable ability to balance content fidelity with style richness, achieving a harmonious integration of both. It ensures that the key structural and motion details of the original content are retained, while preserving the primary style of the object with great precision.
We have also compared our method with existing commercial models, namely Gen-1 [51] and Gen-3, as illustrated in Fig.  6. Our approach demonstrates clear superiority in two critical aspects: achieving a closer style similarity to the reference image and maintaining higher content fidelity to the original video. Moreover, it uniquely excels in preserving the style consistency of local objects, ensuring a more cohesive and visually appealing result, even in complex or dynamic scenes.
5.4 Ablation Studies
We have developed five distinct variants to assess the contribution of its individual components: (1) The full UniVST model, (2) UniVST without mask guidance, (3) UniVST without latent-shift, (4) UniVST without attention-shift, and (5) UniVST without window smoothing. The quantitative outcomes are presented in Table II, while the qualitative results are displayed in Fig. 8.
Our mask-guided strategy effectively maintains the foreground style throughout the transfer process, showing its efficacy both quantitatively and qualitatively, without necessitating additional model components. The latent-shift and attention-shift strategies are both integral to the style transfer process. It is noteworthy noteworthy that the absence of the attention-shift strategy results in an incomplete transfer and a blurring of local details, as exemplified in Fig. 8. Finally, our optical flow smoothing strategy significantly enhances the modeling of temporal consistency, effectively mitigating video artifacts. In addition, we conducted an in-depth study of the three proposed modules on the DAVIS-Laion dataset to further substantiate their individual contributions and overall impact.
Figure 7: Visual results of mask propagation on the DAVIS dataset [42]. The mask is blended with the original image to emphasize the subject, highlighting it while dimming the background.
Figure 8: Visual results for ablations. Various components in our framework have all played important roles.
Figure 9: Quantitative results of mask propagation with different sampling ratio $r$ on the DAVIS [42] dataset. Best view with zooming in.
TABLE III: Performance and efficiency of mask propagation with different anchor frames. The bold fonts represent our final configuration.
[…]
Previous-1(Naive) 0.5321 0.6505 2.365
Previous-1+First 0.7799 0.8714 2.423($\downarrow$2.5%)
Previous-3+First 0.7792 0.8711 2.477($\downarrow$4.7%)
Previous-5+First 0.7800 0.8716 2.577($\downarrow$9.0%)
Previous-7+First 0.7809 0.8721 2.673($\downarrow$13.0%)
Previous-9+First 0.7816 0.8726 2.742($\downarrow$15.9%)
Previous-11+First 0.7820 0.8728 2.868($\downarrow$21.3%)
Previous-13+First 0.7820 0.8728 2.873($\downarrow$21.5%)
Previous-15+First 0.7820 0.8729 2.878($\downarrow$21.7%)
TABLE IV: Performance and efficiency of mask propagation with different values of $k$. The bold fonts represent our final configuration.
[…]
40 0.7806 0.8719 2.467
30 0.7836 0.8739 2.458
20 0.7836 0.8750 2.455
15 0.7885 0.8772 2.452
10 0.7875 0.8766 2.452
Mask Propagation. We first present the visualization of the final mask propagation results, as shown in Fig. 7. Then, we further discuss the impact of downsampling rates on propagation accuracy and inference speed using the DAVIS dataset [42]. Detailed results can be found in Fig. 9. As the sampling rate decreases, the efficiency of our mask propagation is significantly improved, while effectively maintaining the accuracy of the final propagation results. We also examine the impact of anchor frames selection on propagation accuracy and efficiency. As shown in Table III, introducing more anchor frames improves propagation accuracy but also results in longer propagation times. Furthermore, we find that as the number of frames increases, the accuracy gains become less significant. After weighing these trade-offs, we select the first frame and the previous 9 frames as the final hyperparameter configuration. In addition, we perform an ablation study on the hyperparameter $k$ of $k$-NN [33]. As shown in Table IV, the highest propagation accuracy is achieved at $k=15$, without affecting propagation efficiency.
TABLE V: Quantitative results of attention-shift with different timestep intervals on the DAVIS-Laion dataset. The bold fonts represent our final configuration.
Timesteps $\leq 0.6T$ $\leq 0.5T$ $\leq 0.4T$ $\leq 0.3T$ $\leq 0.2T$ $\geq\mathbf{0.4T}$ $\geq 0.5T$ $\geq 0.6T$ $\geq 0.7T$ $\geq 0.8T$
[…]
TABLE VI: Quantitative results of latent-shift with different timestep intervals on the DAVIS-Laion dataset. The bold fonts show our final configuration.
[…]
TABLE VII: Comparison between attention-shift and some existing training-free stylization methods on the DAVIS-Laion dataset. The bold fonts show our final configuration.
Configuration The key-value replacement The key-value AdaIN Ours with Increasing Ours with Decreasing
[…]
Figure 10: Visual results of latent-shift with different timestep intervals on the DAVIS-Laion.
Figure 11: Visual results of attention-shift with different timestep intervals on the DAVIS-Laion. Best view with zooming in.
AdaIN-Guided Style Transfer. We investigate the optimal timestep interval for this method in Table V. It means that applying attention-shift in the early and middle stages of denoising can yield better results. Futhermore, we explore the optimal timestep intervals for latent-shift, with quantitative results provided in Table VI. Applying latent-shift in the late stages yields better results. In addition, we present visual results of these ablation experiments, as shown in Fig. 10 and Fig. 11. For latent-shift, applying it in the early stages of denoising can cause issues such as content blurriness and semantic loss during the transfer process. In contrast, attention-shift produces better visual results when applied in the early stages of denoising than in the later stages. Finally, to further investigate the effectiveness of our proposed attention-shift and the monotonicity of the time coefficient $\beta_{t}$, we compare the performance of four configurations: key-value replacement, key-value AdaIN, ours with increasing time coefficients, and ours with decreasing time coefficients. As shown in Table VII, using a decreasing time coefficient enables attention-shift to outperform other training-free methods.
Figure 12: Qualitative comparisons of different inversion methods on transfer quality. EasyInv [34] demonstrates better transfer quality compared to DDIM-Inv [16].
Figure 13: Visual results of sliding-window consistent smoothing with different window sizes. Best view with zooming in.
Sliding-Window Consistency Smoothing. To assess the impact of sliding window size on editing results and computational costs, we present the visual effects and corresponding inference times for different window sizes in Fig. 13. While larger window sizes yield slight visual improvements, they also result in significant inference costs. On the other hand, we further explore how applying this strategy at different timestep intervals affects final temporal consistency. As shown in Fig. 14, we find that the quality of smoothing strategy varies across the early, middle, and late stages of the denoising process, with the best results achieved when it is applied during the middle stage.
Figure 14: Quantitative results of sliding-window consistent smoothing with varying timestep intervals on the DAVIS-Laion.
Impact of Inversion on UniVST. As a training-free stylization method, we need to obtain the initial noise of the original video, meaning that our method can not bypass inversion technique. Therefore, the performance of this method largely depends on the quality of the initial noise obtained through inversion. In Fig. 12, we find that naive inversion can lead to a loss of semantic details during the transfer process and adversely affect subsequent mask propagation. To minimize the errors introduced by the inversion process, we adopt EasyInv [34] and show better visual effects.
Figure 15: Two examples of multi-step style transfer are demonstrated: the first involves stylizing the background, followed by an additional stylization of the foreground, building upon the initial transformation.
6 Multi-Step Video Style Transfer
Apart from the cases presented in the experimental section, UniVST showcases its versatility by enabling multi-step video style transfer, which allows for more customized transformations and creative potential of video editing. For instance, users can start by stylizing the background of the video using a mask to isolate it from the rest of the scene. After completing this step, they can invert the mask to focus on the previously excluded foreground and apply a separate style transfer to it. This sequential process not only provides greater control but also enables creative layering of styles. As illustrated in Fig. 15, employing distinct style images for the background and foreground in these two stages produces a visually striking and dynamic effect, enhancing the artistic flexibility of the method.
7 Limitation and Future Work
While our UniVST framework shows a clear advantage in localized video style transformation, it has limitations: (1) The accuracy of mask propagation and video content fidelity may be affected by inaccuracies in the inversion process. (2) The style transfer mechanism, guided solely by prior knowledge of LDM [29], may struggle with certain style images, especially those with nuanced color distributions, leading to suboptimal performance. (3) The sliding window smoothing technique, based on optical flow, may be less effective with fast-moving objects, resulting in unsatisfactory outcomes. Future research could focus on developing improved transfer techniques that may enhance visual effects and better manage scenes with rapid motion.
8 Conclusion
This paper presents UniVST, a unified framework for localized video style transfer that features a training-free approach, marking a significant advancement over traditional methods that require full video style transfer. Our contributions include: (1) A point-matching mask propagation strategy that eliminates the need for tracking models, enabling streamlined style transfer to specific video objects. (2) A training-free AdaIN-guided video style transfer mechanism that operates at both the latent and attention levels, ensuring a balance between content fidelity and style richness. (3) A sliding-window consistent smoothing scheme that uses optical flow to refine noise and update the latent space, improving temporal consistency and reducing artifacts. Extensive experiments show that UniVST outperforms existing methods in both qualitative and quantitative assessments, preserving the primary object’s style while ensuring temporal consistency and detail preservation. Futhermore, it can be extended to a wider range of customized models.