### v4

# UniVST: A Unified Framework for Training-free Localized Video Style Transfer

Abstract
This paper presents UniVST, a unified framework for localized video style transfer based on diffusion models. It operates without the need for training, offering a distinct advantage over existing diffusion methods that transfer style across entire videos. The endeavors of this paper comprise: (1) A point-matching mask propagation strategy that leverages the feature maps from the DDIM inversion. This streamlines the model’s architecture by obviating the need for tracking models. (2) A training-free AdaIN-guided localized video stylization mechanism that operates at both the latent and attention levels. This balances content fidelity and style richness, mitigating the loss of localized details commonly associated with direct video stylization. (3) A sliding-window consistent smoothing scheme that harnesses optical flow within the pixel representation and refines predicted noise to update the latent space. This significantly enhances temporal consistency and diminishes artifacts in stylized video. Our proposed UniVST has been validated to be superior to existing methods in quantitative and qualitative metrics. It adeptly addresses the challenges of preserving the primary object’s style while ensuring temporal consistency and detail preservation. Our code is available at https://github.com/QuanjianSong/UniVST.
Index Terms:
Localized Video Style Transfer, Diffusion Models, Training-free.
1 Introduction
Video editing has greatly improved thanks to the use of diffusion models [1, 2, 3]. T2V-Zero [4] changes self-attention mechanisms to cross-frame attention. Tune-A-Video customizes video editing by adjusting attention weights [5]. Fate-Zero [6] keeps the video content intact via information from the inversion process. Animate-Zero [7] and VideoBooth [8] add layers that stress time.
Figure 1: Existing methods suffer from (a) lack of fine-grained control, (b) imbalance between content fidelity and style richness, and (c) temporal inconsistency, whereas our UniVST effectively overcomes these challenges. The last row visualizes the coherence across frames using MimicMotion’s Y-T slices [9] for both BIVDiff [10] and our UniVST.
Video stylization, a sub-area of video editing, is gaining popularity. It is about adding artistic styles to video content. Our review identifies two main video stylization methods. The first method uses the image or text to direct the style. StyleCrafter [11] personalizes video style with an extra training adapter; VideoBooth [8] uses the image for custom video creation prompts; while Diffutoon [12] focuses on cartoon coloring from a text prompt. The second method includes general video editing frameworks like AnyV2V [13] and BIVDiff [10] that can perform various tasks, including style transfer. These usually apply a consistent style across the whole video, which is good for thematic expression but sometimes lacking in precision for specific uses. For example, as shown in Fig. 1, in film, scenes might be shot normally and later stylized to match different environments. This enables varied scene changes from one video take, which is also useful in advertising for efficiently applying local style changes to existing videos. Therefore, selectively styling parts of a video while keeping others unchanged is an important academic and commercial task.
We illustrate Fig. 1 and pinpoint three major limitations of current methods: 1. Lack of Fine-grained Control. While some existing approaches [6, 5] leverage localized text descriptions to preserve the overall style transfer and maintain the primary object’s style, the control is usually too coarse. This lack of precision can lead to the model failing to understand the nuances, potentially causing unintended style transfer effects on objects not explicitly mentioned in the text. 2. Content Fidelity vs. Style Richness. In the realm of video stylization, there is a delicate balance to be struck between content fidelity and style richness. An overemphasis on style richness can result in blurring and alter the original layout of the video, whereas an excessive focus on content fidelity might yield a stylized outcome that is indistinguishable from the original. 3. Temporal Inconsistency. Unlike image stylization, video stylization requires careful consideration of temporal coherence between frames. As in BIVDiff [10] and AnyV2V [13], simply extending image stylization techniques to video often results in frame inconsistencies, causing various flickering and artifacts.
In this paper, we propose UniVST, a unified localized [^1] Here, “localized” indicates a particular part of the video. video style transfer framework that addresses the above issues in a training-free manner. The first step in achieving this is to obtain masks of the region specified by users, which allows for precise control during the subsequent stylization process. However, unlike in the image domain [14, 15], where masks can be readily obtained using additional models, the video domain incurs additional inference overhead and presents a more labor-intensive challenge due to the need for frame-by-frame mask generation. Recognizing the high redundancy between video frames, we propose a Point-Matching Mask Propagation strategy to efficiently extend the initial frame’s mask to subsequent frames. Leveraging feature maps from the inversion process [16], this strategy captures correlations in the latent space, eliminating the need for additional models.
By applying the inversion on both the original video and style image, we obtain the initial noise for each. Integrated with our point-matching mask propagation strategy, we can also obtain per-frame masks for the object region specified by users. Then, we establish new edited latents and construct a three-branch denoising architecture with the content, edited, and style branches, for training-free AdaIN-Guided Localized Video Stylization. It first utilizes localized latent blending to fuse information from the content branch and edited branch, achieving localized style control during the style transfer process. In addition, to enhance information interaction between edited and style branches during denoising, we propose a training-free AdaIN-guided latent-shift and attention-shift mechanism. Unlike the image stylization method [17], It effectively mitigates detail loss while adeptly balancing content fidelity and style richness during the stylization.
However, directly applying AdaIN-guided localized video stylization per-frame will introduce flickering and artifacts in final edited video. We empirically observe that this issue arises from a training-free stylization mechanism, which disrupts the temporal consistency of the edited latent. Inspired by previous works [2, 1], we further employ optical flow to enhance temporal consistency in the latent space. Specifically, we present a specialized Sliding-Window Consistent Smoothing scheme during denoising, which employs optical flow for spatial alignment at the pixel level and refines predicted noise in the latent space. As illustrated in Fig. 1, our smoothing scheme reduces flickering and artifacts while enhancing the temporal consistency in the final edited video.
Overall, our major contributions in this paper are as follows:
  • We introduce, to the best of our knowledge, the first localized video stylization framework with diffusion models, featuring a novel point-matching mask propagation strategy.
  • We develop a training-free AdaIN-guided localized video stylization mechanism that operates at both the latent and attention levels. It effectively harmonizes content fidelity and style richness during the style transfer process.
  • We present a sliding-window consistent smoothing scheme based on optical flow, which adeptly upholds temporal consistency throughout the style transfer process.
  • Comprehensive experiments demonstrate that our framework outperforms several state-of-the-art methods across various qualitative and quantitative evaluation metrics.
2 Related Work
Image Style Transfer. Image style transfer aims to endow an image with a specific style, with early research focusing on non-diffusion methods. RSCT [18] calls for the separate transfer of foreground and background styles, combining them to ensure consistent styling. CAST [19] uses contrastive learning to extract style representations from image features, enhancing performance in arbitrary style transfer tasks. EFDM [20] utilizes exact feature distribution matching to align complex feature distributions across diverse style domains. With the development of diffusion models, recent studies have increasingly explored their application in style transfer. DiffStyle [21] employs LoRA [22] to refine image prompts for stylization, while ArtBank [23] enhances text embeddings through fine-tuning image prompts. StyleID [17] leverages key-value replacement coupled with an initial latent AdaIN [24] to preserve content fidelity. Z* [25] integrates content with diverse styles through an attention-based rearrangement mechanism. InstanceStylePlus [26] utilizes adapters to safeguard content retention and enhance stylistic expression. Although these methods demonstrate superior performance, directly applying them to video tasks using frame-wise processing often fails to ensure semantic continuity.
Video Style Transfer. Video style transfer prioritizes per-frame image style transfer while ensuring visual coherence. Similar to image style transfer, early research also concentrated on non-diffusion approaches. RT-NVST [27] uses feedforward networks for real-time video style transfer with a hybrid loss ensuring temporal consistency. RT-LVST [28] proposes a deep neural network for transferring artistic styles to specific video regions while preserving photorealism. These traditional non-diffusion models incur extra training costs and underperform diffusion models. Recent studies have increasingly employed diffusion models to accomplish video stylization. OCD [29] integrates object-centric sampling and merging, expediting the editing process. StyleCrafter [11] refines adapters using image prompts based on the T2V model. Style-A-Video [30] enhances stylistic alignment through strategic style guidance. Owing to the substantial computational demands of training, some studies have turned to training-free video stylization. Diffutoon [12] pioneers primary and editing branches for multi-segment editing in video colorization. BIVDiff [10] combines frame-by-frame image editing with temporal consistency modeling. AnyV2V [13] follows a similar initial editing strategy, subsequently leveraging an I2V model [31] to propagate changes across the video sequence. Although these diffusion-based methods exhibit strong performance, they struggle with fine-grained local editing while balancing content fidelity and style richness. In contrast, our proposed UniVST achieves fine-grained transfer while maintaining both content fidelity and style richness in a training-free manner.
Figure 2: Overall framework of our UniVST. It consists of three core components. (1) Point-Matching Mask Propagation, (2) AdaIN-Guided Localized Video Stylization (Localized Latent Blending, AdaIN-Guided Latent-Shift and Attention-Shift) and (3) Sliding-Window Consistent Smoothing.
3 Preliminary
Latent Diffusion Model. LDM [32] adds and removes noise in a low-dim space by using the encoder $\mathcal{E}$ and decoder $\mathcal{D}$. Given a Gaussian noise $Z_{T}$, DDIM denoising [3] can be formulated as:
$Z_{t-1}=\sqrt{\alpha_{t-1}}\underbrace{{Z}_{t\rightarrow 0}}_{\text{predicted }Z_{0}}+\underbrace{\sqrt{1-\alpha_{t-1}}\epsilon_{\theta}\left(Z_{t},t,C\right)}_{\text{direction pointing to }Z_{t-1}},$ (1)
in which ${Z}_{t\rightarrow 0}$ is an approximate estimation of $Z_{0}$ at time step $t$:
${Z}_{t\rightarrow 0}=\left(Z_{t}-\sqrt{1-{\alpha}_{t}}\epsilon_{\theta}\left(Z_{t},t,C\right)\right)/\sqrt{\alpha_{t}},$ (2)
where $\alpha_{t}$ is predefined parameter and $C=\phi$ for editing integrity.
DDIM Inversion. DDIM [16] transforms noise $Z_{T}$ into $Z_{0}$. Given reversible ODE [33], its reverse process can be described by:
${Z}_{t+1}={A}_{t}{Z}_{t}+{B}_{t}\epsilon_{\theta}\left({Z}_{t},t,\phi\right),$ (3)
where $A_{t}$ and $B_{t}$ are the functions of time step $t$. During editing, the initial noise $Z_{T}$, which closely aligns with the original video distribution $Z_{0}$, can be accurately derived using this equation.
3D UNet Extension. Following previous works [34, 1], standard LDM will be extended into a 3D U-Net for video tasks. The $3\times 3$ convolutional kernels in convolutional blocks are expanded to $1\times 3\times 3$. To enable frame parallelization, the intermediate features are reshaped from $\mathbb{R}^{b\times c\times f\times h\times w}$ to $\mathbb{R}^{(bf)\times c\times h\times w}$ before self-attention and cross-attention. To enhance frame interaction, we follow T2V-Zero [4] to modify the self-attention into cross-frame attention:
$\begin{array}[]{c}Q^{i}=W^{query}\cdot\mathcal{F}^{i},\quad K^{i}=W^{key}\cdot[\mathcal{F}^{{1}},\mathcal{F}^{{i-1}}],\\ V^{i}=W^{value}\cdot[\mathcal{F}^{{1}},\mathcal{F}^{{i-1}}],\end{array}$ (4)
where $\mathcal{F}^{i}$ denotes the extracted feature of the $i$-th frame, while $W^{query}$, $W^{key}$, and $W^{value}$ represent the corresponding learnable mapping matrices, with [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] indicating concatenation operation.
4 Localized Video Style Transfer
Given an original video with $N$ frames $\{I^{i}\}_{i=1}^{N}$ and a style image $I^{s}$, our objective is to transform $\{I^{i}\}_{i=1}^{N}$ into a new sequence $\{\mathcal{I}^{i}\}_{i=1}^{N}$, which incorporates the reference style while preserving the primary objects’ original style. To achieve this, we propose UniVST, a unified framework for training-free localized video style transfer. The overall framework is illustrated in Fig. 2. UniVST first applies DDIM inversion to the original video and style image to obtain their initial noise and integrates Point-Matching Mask Propagation to generate masks for the object regions. It then performs AdaIN-Guided Localized Video Stylization with a three-branch architecture for information interaction. Moreover, Sliding-Window Consistent Smoothing is incorporated into the denoising process, enhancing the temporal consistency in the latent space.
Figure 3: The comparison between our mask propagation strategy and the traditional SAM model in terms of memory usage and inference time.
Algorithm 1 Efficient Point-Matching Mask Propagation
1:Input: $\mathcal{M}_{1}$ The mask for the first frame.
2:Input: $F_{t_{0}}$ The feature map from last up-sampling block.
3:Input: $r$ The sampling rates.
4:Input: $k$ The $k$-nearest points.
5:Input: $n$ The number of reference frames.
6:Output: $\{\mathcal{M}^{i}\}_{i=1}^{N}$ The masks for all frames.
7:Initialize: PreviousFeatures [$\leftarrow$[\,]← [ ] Initialize empty list for previous features.
8:Initialize: PreviousMasks [$\leftarrow$[\,]← [ ] Initialize empty list for previous masks.
9:Initialize: FirstFeature [$\leftarrow$[\,]← [ ] Initialize empty list for first feature.
10:for $i=1,\dots,N$ do
11:    $\text{ForeIndices}\leftarrow\text{Where}(M_{i-1}=1)$
12:    $\text{ForeIndices}\leftarrow\text{Where}(M_{i-1}=0)$
13:    $\text{ForeIndices}\leftarrow\operatorname{RandomSample}(\text{ForeIndices},r\cdot\frac{\lvert\text{ForeIndices}\rvert}{\lvert\text{ForeIndices}\rvert+\lvert\text{BackIndices}\rvert})$
14:    $\text{BackIndices}\leftarrow\operatorname{RandomSample}(\text{BackIndices},r\cdot\frac{\lvert\text{BackIndices}\rvert}{\lvert\text{ForeIndices}\rvert+\lvert\text{BackIndices}\rvert})$
15:    $\text{CurrentIndices}\leftarrow\operatorname{Concat}(\text{ForeIndices},\text{BackIndices})$
16:    $\text{CurrentFeature}\leftarrow F_{t_{0}}^{i}[\text{CurrentIndices}]$
17:    if $i=1$ then
18:       FirstFeature.append(CurrentFeature) Store the first frame.
19:       $\mathcal{M}^{i}\leftarrow\mathcal{M}^{1}$
20:    else
[…]
22:          PreviousFeatures.pop(0) Remove the oldest frame feature.
23:          PreviousMasks.pop(0) Remove the oldest frame mask.
24:       end if
[…]
26:       $\mathcal{M}^{i}\leftarrow\text{KNN}(f,F_{t_{0}}^{i},\operatorname{Concat}(\text{PreviousMasks},\mathcal{M}^{1}),k)$
27:       PreviousFeatures.append(CurrentFeature) Store the current frame.
28:       PreviousMasks.append($\mathcal{M}_{i}$) Store the current mask.
29:    end if
30:end for
31:Return: $\{\mathcal{M}^{i}\}_{i=1}^{N}$
4.1 Point-Matching Mask Propagation
“A towering building begins with a solid foundation.”
As the proverb clearly illustrates, the foundation is crucial for any significant endeavor. In the context of the localized video style transfer, the first essential step is to obtain precise and accurate masks for the region of interest specified by the users. These masks essentially form the cornerstone for enabling fine-grained control over specific regions during the subsequent stylization process.
Compared to image style transfer, the video domain presents a unique challenge: obtaining mask information for each frame. One potential solution is to employ segmentation models, such as SAM [35], which can perform frame-by-frame segmentation based on the specified object. However, as shown in Fig. 3, this approach has notable drawbacks. Specifically, it introduces an additional memory overhead of 7.9 GB and is computationally intensive, which is unfavorable for deploying our UniVST framework.
Given the substantial redundancy between consecutive video frames, we propose leveraging this redundancy to efficiently propagate the initial frame’s mask across all subsequent frames, thereby eliminating the need for additional segmentation models. This approach not only reduces memory usage but also accelerates the overall processing time. Thus, we introduce our point-matching mask propagation strategy, which is outlined in Algorithm 1.
Point-Matching. Inspired by DIFT [36], the three up-sampling blocks in the U-Net are replete with rich semantic information. Therefore, during the DDIM inversion in the original video, we can retain the feature map $\{F_{t_{0}}^{i}\}_{i=1}^{N}$ from the last up-sampling block at a given step $t_{0}$, and use them to calculate the pixel correspondences across frames. Specifically, to pinpoint the pixel in the $j$-th frame that most closely resembles a given point $p_{i}$ in the $i$-th frame, we calculate the cosine similarity between the corresponding point $p^{i}$ in the feature map $F_{t_{0}}^{i}$ and all points $p^{j}$ in the feature map $F_{t_{0}}^{j}$:
$p^{j}={\underset{p^{j}\in F_{t_{0}}^{j}}{\arg\min}}\ \operatorname{CosSim}\big{(}p^{i},p^{j}\big{)},\quad p^{i}\in F_{t_{0}}^{i}.$ (5)
In which, we can extend the initial frame’s mask across subsequent frames by the mask propagation strategy introduced below.
Mask Propagation. Given an initial object mask $\mathcal{M}^{1}$ specified by the users, we use $k$-NN [37] to propagate this mask to subsequent frames. For each point in the $i$-th frame, we identify the $k$ most similar points in the $(i-1)$-th frame based on the point-matching principle described above. If the majority of these $k$ points are in the foreground, the corresponding point in the $i$-th frame is classified as the foreground; otherwise, it is classified as the background. In this way, we can obtain the masks $\{\mathcal{M}^{i}\}_{i=1}^{N}$ for all frames.
However, as shown in Fig. 4, using a naive autoregressive approach for mask propagation will result in error accumulation, which degrades the accuracy of subsequent frames. To mitigate this issue, we design the anchor frames mechanism that incorporates information from the first frame as well as the previous $n$ frames. On the other hand, although the introduction of the anchor frames mechanism improves the propagation accuracy of subsequent frames, it introduces additional similarity calculations, potentially reducing efficiency. To address this issue, we apply random down-sampling at a rate of $r$ to the foreground and background of the anchor frames, with the rate adjusted based on their respective proportions. As illustrated in Fig. 4, the combination of anchor frames and random down-sampling strategies effectively preserves mask propagation accuracy while improving efficiency. More comprehensive analysis will be provided in Sec. 5.4.2.
Figure 4: Comparison of accuracy and inference time under different mask propagation strategies. The accuracy is measured by relevant segmentation metrics: the intersection over union (IoU) and dice coefficients (Dice), which assess the overlap between predicted and ground truth regions. Our strategy effectively balances inference time and accuracy.
4.2 AdaIN-Guided Localized Video Stylization
To achieve localized video stylization, we first separately obtain the initial latent of the original video $\{I^{i}\}_{i=1}^{N}$ and the style image $I^{s}$ using the inversion technique [38], denoted as $Z_{t}$ and $Z_{t}^{s}$ respectively, for $t=1\rightarrow T$. Integrated with our point-matching mask propagation strategy, we can also obtain the mask frame-by-frame efficiently. We then initialize the edited latent $\mathcal{Z}_{T}=Z_{T}$ and introduce a three-branch architecture with content, edited, and style branches, which perform the subsequent denoising process in parallel for $t=T\to 1$. These branches will interact through localized latent blending, AdaIN-guided latent-shift, and AdaIN-guided attention-shift, enabling precise control in subsequent video stylization.
Localized Latent Blending. To ensure that the style within the user-specified region is preserved throughout the denoising process, we utilize the masks $\{\mathcal{M}^{i}\}_{i=1}^{N}$, which is obtained from the previous mask propagation, as auxiliary constraints. Specifically, during each denoising step, we blend the latent of content branch $Z_{t}^{i}$ and the latent of style branch $\mathcal{Z}_{t}^{i}$, using the mask $\mathcal{M}^{i}$ of the $i$-th frame:
$\mathcal{Z}_{t}^{i}=\mathcal{M}^{i}\cdot Z_{t}^{i}+(1-\mathcal{M}^{i})\cdot\mathcal{Z}_{t}^{i},\quad i=1,...,N$ (6)
This blending operation effectively ensures that the style within the user-specified region is seamlessly retained while allowing the rest of the frame to be fully stylized according to the edited branch.
AdaIN-Guided Latent-Shift. AdaIN [24] is primarily used in image processing [17], where it adjusts the mean and standard deviation of an image to control its color style. Inspired by this, we apply the AdaIN technique to integrate the color style of the reference image into the edited video. Specifically, within the interval $t\in[\tau_{0},\tau_{1}]$, we apply AdaIN operation to the latent $\mathcal{Z}_{t}$ of the editing branch, shifting its mean and standard deviation to match those in the latent $Z_{t}^{s}$ of the style branch, as follows:
$\tilde{\mathcal{Z}}_{t}$ $=\text{AdaIN}(\mathcal{Z}_{t},Z_{t}^{s}),$ (7)
$\text{AdaIN}(\mathcal{Z}_{t},Z_{t}^{s})=\sigma\left(Z_{t}^{s}\right)\left(\frac{\mathcal{Z}_{t}-\mu\left(\mathcal{Z}_{t}\right)}{\sigma\left(\mathcal{Z}_{t}\right)}\right)+\mu\left(Z_{t}^{s}\right),$ (8)
where $\mu(\cdot)$ and $\sigma(\cdot)$ denote respective mean and standard deviation.
Figure 5: Comparison of training-free stylization methods: (a) original video frame, (b) key-value replacement [17], (c) key-value AdaIN and (d) AdaIN-guided attention-shift. Our approach mitigates local detail loss.
AdaIN-Guided Attention-Shift. In addition to transferring color tone, we also need to incorporate the overall style of the reference image into the edited video. The previous work StyleID [17] achieves style transfer by modifying the standard self-attention process. Specifically, the query $\mathcal{Q}_{t}$ in the editing branch attends to the key $K_{t}^{s}$ and value $V_{t}^{s}$ of the style branch instead of its own, referred to as K-V replacement. Furthermore, to eliminate content damage caused by direct replacement, [17] blends the original video’s query $Q_{t}$ with the edited query $\mathcal{Q}_{t}$. It is formulated as:
$\begin{split}\tilde{\mathcal{Q}}_{t}&=\gamma\cdot\mathcal{Q}_{t}+(1-\gamma)\cdot Q_{t},\\ &\tilde{\mathcal{K}}_{t}=K_{t}^{s},\;\tilde{\mathcal{V}}_{t}=V_{t}^{s},\\ &\operatorname{Attention}(\tilde{\mathcal{Q}}_{t},\tilde{\mathcal{K}}_{t},\tilde{\mathcal{V}}_{t}),\end{split}$ (9)
where $\gamma$ controls the balance between content and style. However, as observed in Fig. 5(b), K-V replacement still results in slight local artifacts in the stylized videos. To further reduce the local artifacts caused by direct replacement, we apply AdaIN to $\mathcal{K}_{t}$ and $\mathcal{V}_{t}$ in the editing branch, shifting their mean and standard deviation to align with $K^{s}$ and $V^{s}$ in the style branch, which can be described as:
$\tilde{\mathcal{K}}_{t}=\text{AdaIN}(\mathcal{K}_{t},K_{t}^{s}),\quad\tilde{\mathcal{V}}_{t}=\text{AdaIN}(\mathcal{V}_{t},V_{t}^{s}).$ (10)
Though the K-V AdaIN alleviates the artifact phenomenon, as indicated in Fig. 5(c), it introduces new local detail blurring. Upon analysis, we find that early denoising stages have less semantic information, making them suitable for K-V AdaIN, while later stages, containing more semantic information, are suited for key-value replacement. In this way, within the interval $t\in[\tau_{2},\tau_{3}]$, we gradually decrease the impact of the K-V AdaIN by combining the K-V replacements with a decreasing weight $\beta_{t}$ as follows:
$\tilde{\mathcal{K}}_{t}$ $=\beta_{t}\cdot\text{AdaIN}(\mathcal{K}_{t},K_{t}^{s})+(1-\beta_{t})\cdot K_{t}^{s},$ (11)
$\tilde{\mathcal{V}}_{t}$ $=\beta_{t}\cdot\text{AdaIN}(\mathcal{V}_{t},V_{t}^{s})+(1-\beta_{t})\cdot V_{t}^{s},$
$\beta_{t}$ $=\frac{\beta_{\tau_{3}}-\beta_{\tau_{2}}}{\tau_{3}-\tau_{2}}\cdot(t-\tau_{2})+\beta_{\tau_{2}},$
in which, $\beta_{\tau_{2}}$ and $\beta_{\tau_{3}}$ are the predefined weighting factors.
By dynamically combining K-V AdaIN with K-V replacement in Eq. (11) and query blending in Eq. (9), as shown in Fig. 5(d), the stylization results achieve an effective balance between content fidelity and style richness. Further quantitative comparison with other stylization methods will typically be presented in Sec. 5.4.3.
4.3 Sliding-Window Consistent Smoothing
Flickering and artifacts in the final edited video occur when AdaIN-guided localized video stylization is applied frame by frame. We empirically observe that this issue is caused by the training-free stylization mechanism, which disrupts the temporal consistency of the edited latent. Inspired by previous works [1, 2], we further incorporate optical flow with specialized sliding-window consistent smoothing to improve inter-frame coherence during denoising.
Optical Flow Warping. Optical flow estimation is a pivotal technique widely employed in object tracking [39, 40]. It estimates pixel-wise horizontal and vertical motion velocities by analyzing pixel variations across frames. Building on this, we utilize the optical flow map to align semantic content across frames, enhancing the coherence of the edited video. Specifically, given two images, $A$ and $B$, we first employ RAFT [41] to estimate the optical flow map from $A$ to $B$. This map is then used to transform $B$, aligning it with $A$ to ensure precise spatial correspondence. In cases of occlusion, bidirectional optical flow can identify occluded regions and recover them with relevant content from $A$. The above process, termed $\operatorname{Warp}(A,B)$, is essential for the following discussions.
Sliding-Window Smoothing. Rather than relying on simple post‐processing, we integrate the Warp operation into the denoising process to achieve local smoothing. At time step $t$, we calculate the latent representation $\mathcal{Z}_{t\rightarrow 0}$ from the predicted noise $\varepsilon_{t}$ according to Eq. (2). $\mathcal{Z}_{t\rightarrow 0}$ is then decoded by the decoder $\mathcal{D}$ into a pixel representation $\{P_{t}^{i}\}_{i=1}^{N}$. For each decoded frame $P_{t}^{i}$, we apply a window of size $2m$ to warp neighboring frames onto it, followed by averaging to smooth the original pixel of $P_{t}^{i}$, as detailed below:
$\bar{P}_{t}^{i}\leftarrow\frac{1}{2m+1}\cdot\sum_{j=i-m}^{i+m}\operatorname{Warp}(P_{t}^{i},P_{t}^{j}).$ (12)
The smoothed frame from the current window is used in the next window’s calculations, as shown in Fig. 2. Next, the VAE’s encoder $\mathcal{E}$ re-encodes the smoothed $\{\bar{P}_{t}^{i}\}_{i=1}^{N}$ into latent representations $\bar{\mathcal{Z}}_{t\rightarrow 0}$, which is a refined version of $\mathcal{Z}_{t\rightarrow 0}$. Subsequently, the predicted noise $\varepsilon_{t}$ is adjusted using the inversion of Eq. (2):
$\bar{\varepsilon}_{t}\leftarrow\frac{\mathcal{Z}_{t}-\sqrt{\alpha_{t}}\bar{\mathcal{Z}}_{t\rightarrow 0}}{\sqrt{1-\alpha_{t}}}.$ (13)
Finally, the latent representation is updated according to the DDIM schedule in Eq. (1), enhancing coherence across frames:
$\mathcal{Z}_{t-1}\leftarrow\sqrt{\alpha_{t-1}}\bar{\mathcal{Z}}_{t\rightarrow 0}+\sqrt{1-\alpha_{t-1}}\bar{\varepsilon}_{t}.$ (14)
The above sliding-window consistent smoothing scheme is performed only within the interval $t\in[\tau_{4},\tau_{5}]$. Moreover, due to the challenges of optical flow estimation for fast-moving objects, we avoid global smoothing for long-range spatial alignment, opting instead for local sliding-window for short-range spatial alignment.
TABLE I: Quantitative comparison on DAVTG-WikiArt, where * denotes non-diffusion methods. The bold and underline indicate the best and second-best.
Methods Overall transfer performance Local style preservation Temporal consistency Inference efficiency
[…]
Diffutoon [12] 43.150 26.268 0.582 0.961 0.954 0.980 44.263
StyleCrafter [11] 39.119 21.008 0.777 0.928 0.895 0.978 111.225
AnyV2V [13] 33.774 19.104 0.680 0.938 0.914 0.978 78.474
BIVDiff [10] 46.071 25.468 0.740 0.921 0.894 0.955 158.367
EFDM [20] 37.407 22.350 0.602 0.936 0.935 0.963 5.969
CAST [19] 45.198 26.174 0.663 0.948 0.936 0.975 7.348
UniVST (ours) 37.152 26.134 0.369 0.986 0.991 0.980 162.745
TABLE II: Quantitative comparison on DAVTG-Laion, where * denotes non-diffusion methods. The bold and underline indicate the best and second-best.
Methods Overall transfer performance Local style preservation Temporal consistency Inference efficiency
[…]
Diffutoon [12] 38.610 23.700 0.563 0.959 0.953 0.979 44.263
StyleCrafter [11] 43.161 23.246 0.780 0.926 0.897 0.978 111.225
AnyV2V [13] 35.648 20.019 0.696 0.936 0.918 0.975 78.474
BIVDiff [10] 36.676 20.156 0.733 0.921 0.897 0.956 158.367
EFDM [20] 35.673 21.549 0.582 0.933 0.931 0.964 5.969
CAST [19] 34.828 19.974 0.661 0.947 0.936 0.968 7.348
UniVST (ours) 30.636 21.146 0.383 0.986 0.990 0.981 162.745
5 Experimentation
5.1 Experimental Settings
Our proposed UniVST is based on the Stable Diffusion V-1.5 [32]. The hyper-parameters are configured as follows: $\tau_{0}=0.1T$, $\tau_{1}=0.2T$, $\tau_{2}=0.4T$, $\tau_{3}=1.0T$, $\tau_{4}=0.5T$, $\tau_{5}=0.6T$, $t_{0}=0.4T$, with the total time period $T$ set to 50. Additionally, we set $\gamma=0.35$, $\beta_{\tau_{2}}=0.1$, $\beta_{\tau_{3}}=0.9$, $r=0.3$, $k=15$, $m=2$, and $n=9$. Original videos are resized to $512\times 512$ pixels and processed in batches of 16 frames. We use the RAFT model [41] for optical flow estimation during the denoising process. UniVST can be executed on a single NVIDIA GeForce RTX 3090 GPU without additional training or fine-tuning, making it both efficient and practical for deployment.
Datasets Details. Building upon the foundational work referenced in [17, 30], we have meticulously selected style images from two renowned databases: WikiArt [42] and Laion-Aesthetics-6.5+ [43], each contributing a collection of 20 distinct images. For the content videos, we have procured a dataset comprising 50 videos from the DAVIS2016 [44] and TGVE [45] datasets, which we combined and referred to as DAVTG. Through the integration of these datasets, we have successfully established two comprehensive subsets: DAVTG-WikiArt, which contains 1,000 curated pairs, and DAVTG-Laion, which also features 1,000 paired entries.
Baselines. We compare our UniVST with six advanced methods. These include two video stylization techniques: Diffutoon [12] and StyleCrafter [11]; two video editing frameworks: AnyV2V [13] and BIVDiff [10]; and two traditional non-diffusion methods: EFDM [20] and CAST [19]. Since our specific task does not perfectly align with some of the existing methods, we have made the necessary adjustments to ensure a fair comparison. In the case of Diffutoon [12], which is designed to process text inputs for coloring tasks and does not natively support the input of style images. To overcome this limitation, we leverage GPT-4o [46] to generate descriptive captions for the style images. For the style generation task with StyleCrafter [11], which does not inherently support video editing, we employ inversion to obtain the initial noise from the original video, which is then utilized in the generation process. For the unified video editing frameworks, AnyV2V [13] and BIVDiff [10], we have integrated InstantStyle [47] as the image style transfer model to perform the video style transfer task. Regarding the traditional non-diffusion methods, EFDM [20] and CAST [19], which are well-aligned with our task, we have retained them in their original form without any modifications.
5.2 Quantitative Comparison
We evaluate the stylized videos generated by different methods across four aspects, in order to comprehensively assess these methods’ performance. (1) Overall transfer performance. Drawing inspiration from the methodologies outlined in [17, 30], we assess the overall transfer performance using ArtFID [48], a comprehensive metric that combines LPIPS [49] for gauging content fidelity and FID [50] for assessing style richness. (2) Local style preservation. To evaluate the preservation of local style, we utilize SSIM [51] metric for the texture preservation, while CLIP-I score [52] is utilized to assess the preservation of semantic content. During the evaluation, the background is neutralized to ensure accuracy. (3) Temporal consistency. In line with the approaches detailed in [30, 11], we measure the similarity between consecutive frames using the CLIP-F score [52], which serves as a measure of inter-frame coherence in the edited video. (4) Inference efficiency. To assess the inference efficiency, we calculate the aggregate time during inference, thereby assessing their performance in terms of computational speed. Furthermore, we employ six established methods as baselines and benchmark them against our UniVST on two datasets: DAVTG-WikiArt and DAVTG-Laion. Quantitative comparisons are comprehensively shown in Table I and Table II.
Figure 6: Qualitative comparison of our proposed UniVST with existing methods. superscript * indicates non-diffusion methods.
The results demonstrate that our UniVST outperforms all competing methods across three performance metrics. On both the DAVTG-WikiArt and DAVTG-Laion datasets, UniVST consistently surpasses existing methods in local style preservation and temporal consistency, showcasing its superiority in localized video style transfer tasks over existing methods. Regarding overall transfer performance, UniVST achieves the best results on the DAVTG-Laion dataset and secures the second spot on DAVTG-WikiArt. Localized transfer tasks might lead to a marginally higher FID score, slightly affecting overall performance compared to other methods. Nevertheless, these findings underscore UniVST’s proficiency in effectively balancing content fidelity and style richness.
Our inference efficiency does not stand out significantly. This is because we opt not to adopt the temporal attention mechanisms in existing T2V frameworks [11, 30]. Instead, we utilize the sliding-window consistent smoothing scheme, which requires dynamic optical flow estimation during the subsequent denoising process. While this approach improves temporal consistency, it also leads to increased computational overhead. The detailed discussion on the inference efficiency of each module will be provided in Sec. 5.5.
TABLE III: Ablation study on the DAVTG-WikiArt dataset. Bold values indicate the best performance.
Model Variants Overall Transfer Performance Local Style Preservation Temporal Consistency
[…]
W/o all components (baseline) 45.676 31.380 0.411 0.979 0.976 0.981
W/o localized latent blending 38.014 26.050 0.405 0.966 0.970 0.979
W/o AdaIN-guided latent-shift 37.649 26.909 0.349 0.985 0.990 0.980
W/o AdaIN-guided attention-shift 39.589 27.506 0.388 0.985 0.988 0.983
W/o sliding-window consistent smoothing 38.626 26.907 0.384 0.982 0.989 0.970
W/ all components (ours) 37.152 26.134 0.369 0.986 0.991 0.980
5.3 Qualitative Comparison
For qualitative evaluation, we have juxtaposed the editing results with those of six baseline methods, as illustrated in Fig. 6. Diffutoon [12], which is renowned for its expertise in cartoon coloring, exhibits limited responsiveness to a variety of image styles. This results in a final edited video that is notably monotonous, and heavily inclined towards a cartoonish anime aesthetic. StyleCrafter [11], tailored for style-based video generation, demonstrates a high sensitivity to stylistic cues. However, it encounters difficulties in preserving content fidelity, leading to substantial distortion of background details and character features during the transfer process. BIVDiff [10] falls short in both content fidelity and temporal consistency, similarly altering background information and character traits during the transfer. AnyV2V [13] excels in maintaining temporal consistency but faces challenges with transfer tasks involving a central object. It often over-transfers the subject’s style, resulting in a loss of its distinctive attributes. As for the traditional non-diffusion methods, EFDM [20] and CAST [19], they degrade fine details and induce color artifacts during the style transfer process. On the contrary, our UniVST exhibits a remarkable ability to balance content fidelity and style richness, achieving a harmonious integration of both. It ensures that the key structural and motion details of the original videos are retained, while preserving the primary style of the object with great precision.
TABLE IV: Ablation study on the DAVTG-Laion dataset. Bold values indicate the best performance.
Model Variants Overall Transfer Performance Local Style Preservation Temporal Consistency
[…]
W/o all components (baseline) 38.981 26.594 0.413 0.979 0.976 0.981
W/o localized latent blending 30.848 20.974 0.404 0.961 0.967 0.979
W/o AdaIN-guided latent-shift 31.001 21.766 0.362 0.985 0.989 0.980
W/o AdaIN-guided attention-shift 32.534 22.623 0.377 0.985 0.989 0.981
W/o sliding-window consistent smoothing 30.834 21.249 0.386 0.981 0.988 0.971
W/ all components (ours) 30.636 21.146 0.383 0.986 0.990 0.981
Figure 7: Visual results for ablations. Various components in our framework have all played important roles. Best view with zooming-in.
5.4 Ablation Studies
In this section, we first validate the effectiveness of each core component in our proposed UniVST. Then, we conduct an in-depth analysis of hyper-parameter settings for the three modules: point-matching mask propagation, AdaIN-guided localized video stylization, and sliding-window consistency smoothing, individually.
TABLE V: Accuracy and efficiency of mask propagation under different value of $n$ on DAVIS-Laion dataset, Bold fonts represent final configuration.
[…]
n=1 0.5321 0.6505 2.365 (Base)
n=1 + First 0.7799 0.8714 2.423 ($\downarrow$2.5%)
n=3 + First 0.7792 0.8711 2.477 ($\downarrow$4.7%)
n=5 + First 0.7800 0.8716 2.577 ($\downarrow$9.0%)
n=7 + First 0.7809 0.8721 2.673 ($\downarrow$13.0%)
n=9 + First 0.7816 0.8726 2.742 ($\downarrow$15.9%)
n=11 + First 0.7820 0.8728 2.868 ($\downarrow$21.3%)
n=13 + First 0.7820 0.8728 2.873 ($\downarrow$21.5%)
n=15 + First 0.7820 0.8729 2.878 ($\downarrow$21.7%)
Figure 8: Quantitative results of mask propagation with different sampling ratio $r$ on DAVIS [44] dataset. Best view with zooming-in.
TABLE VI: Accuracy and efficiency of mask propagation with different values of $k$ on DAVIS-Laion dataset. Bold fonts represent final configuration.
[…]
40 0.7806 0.8719 2.467
30 0.7836 0.8739 2.458
20 0.7836 0.8750 2.455
15 0.7885 0.8772 2.452
10 0.7875 0.8766 2.452
5.4.1 Ablation Study on Core Components
We have developed six distinct variants to assess the contribution of its individual components: (1) UniVST without all components (baseline), (2) UniVST without localized latent blending, (3) UniVST without AdaIN-guided latent-shift, (4) UniVST without AdaIN-guided attention-shift, (5) UniVST without sliding-window consistent smoothing and (6) UniVST with all components (ours). The quantitative outcomes are presented in Table III and Table IV, while the qualitative results are displayed in Fig. 7. Our mask-guided strategy effectively preserves the foreground style during the stylization process, demonstrating its efficacy quantitatively and qualitatively, without the need for additional models like SAM [35]. The AdaIN-guided latent-shift and attention-shift strategies are essential for style transfer. Notably, omitting the attention-shift strategy leads to incomplete transfer and blurred local details, as shown in Fig. 7. Finally, our sliding-window consistency smoothing scheme enhances temporal consistency and effectively reduces video artifacts. In the following section, we conduct an in-depth study of the three proposed modules on the DAVIS-Laion dataset to further substantiate their individual contributions, respectively.
Figure 9: Visual results of our point-matching mask propagation strategy on the DAVIS dataset [44]. The masks are blended with the original video to emphasize the foreground, while dimming the background. Our propagation strategy can effectively adapt to various complex scenarios.
TABLE VII: Quantitative results of AdaIN-guided attention-shift with different timestep intervals on DAVIS-Laion dataset. Bold fonts indicate final configuration.
Timestep Intervals $\leq 0.6T$ $\leq 0.5T$ $\leq 0.4T$ $\leq 0.3T$ $\leq 0.2T$ $\geq\mathbf{0.4T}$ $\geq 0.5T$ $\geq 0.6T$ $\geq 0.7T$ $\geq 0.8T$
[…]
TABLE VIII: Quantitative results of AdaIN-guided latent-shift with different timestep intervals on DAVIS-Laion dataset. Bold fonts indicate final configuration.
Timestep Intervals $\mathbf{0.1T}\sim\mathbf{0.2T}$ $0.2T\sim 0.3T$ $0.4T\sim 0.5T$ $0.5T\sim 0.6T$ $0.7T\sim 0.8T$ $0.8T\sim 0.9T$
[…]
TABLE IX: Comparison of our method with other training-free stylization methods on DAVIS-Laion dataset. Bold values indicate final configuration.
Method K-V Replacement K-V AdaIN Ours ($\beta_{t}\uparrow$) Ours ($\beta_{t}\downarrow$)
[…]
5.4.2 Analysis of Point-Matching Mask Propagation
We first investigate the effect of down-sampling rates on propagation accuracy and inference time using the DAVIS dataset [44]. Detailed results are presented in Fig. 8. As the down-sampling rate $r$ decreases, the efficiency of our mask propagation improves significantly, while maintaining the accuracy of the final propagation results. Next, we also examine the impact of anchor frame selection on propagation accuracy and efficiency. As shown in Table V, increasing the number of anchor frames improves propagation accuracy but also leads to longer propagation times. Moreover, we observe that as the number of frames increases, the accuracy gains diminish. After considering these trade-offs, we choose the first frame and the previous $n=9$ frames as the final hyper-parameter configuration. We also compare the results for different values of the hyper-parameter $k$ in $k$-NN [37]. As shown in Table VI, the highest propagation accuracy is consistently achieved when $k=15$, without sacrificing the overall propagation efficiency.
We present qualitative examples of our mask propagation results in Fig. 9. In the short-term propagation setting, we evaluate scenarios involving fast-moving foreground objects, deformations, and occlusions. Despite these complexities, our method demonstrates precise object tracking and robust mask alignment, consistently preserving accuracy across diverse conditions. For long-term propagation, our approach continues to deliver strong performance. The integration of anchor frames enables stable and accurate mask propagation over extended temporal distances, mitigating typical degradation issues associated with long sequences. However, certain limitations persist, which will be examined in Sec. 5.5.
Figure 10: Visual results of latent-shift with different timestep intervals on DAVIS-Laion dataset. Best view with zooming-in.
Figure 11: Quantitative results of sliding-window consistent smoothing with varying timestep intervals. Best view with zooming-in.
5.4.3 Analysis of AdaIN-Guided Localized Video Stylization
We investigate the optimal time interval for this method in Table VII. It means that applying AdaIN-guided attention-shift in the early and middle stages of denoising can yield better results. Furthermore, we explore the optimal timestep intervals for AdaIN-guided latent-shift, with quantitative results provided in Table VIII. Applying AdaIN-guided latent-shift in the late stages yields better results. In addition, we present visual results of these ablation experiments, as shown in Fig. 10 and Fig. 12. For AdaIN-guided latent-shift, applying it in the early stages of denoising can cause issues such as content blurriness and semantic loss during the transfer process. In contrast, AdaIN-guided attention-shift produces better visual results when applied in the early stages of denoising than in the later stages.
Figure 12: Visual results of attention-shift with different timestep intervals on DAVIS-Laion dataset. Best view with zooming-in.
Figure 13: Visual results of sliding-window consistent smoothing with different window sizes on DAVIS-Laion dataset. Best view with zooming-in.
Finally, to further investigate the effectiveness of our proposed AdaIN-guided attention-shift mechanism and the impact of the monotonicity of the time coefficient $\beta_{t}$, we compare the performance of four configurations: K-V replacement [17], K-V AdaIN, our method with increasing time coefficients, and our method with decreasing time coefficients. As shown in Table IX, using decreasing time coefficient enables the AdaIN-guided attention-shift to outperform other training-free methods, such as K-V replacement, K-V AdaIN, or attention-shift with increasing time coefficients.
Figure 14: Inference time distribution of different components in UniVST. Sliding-window consistent smoothing exhibits an inference bottleneck.
5.4.4 Analysis of Sliding-Window Consistency Smoothing
To assess the impact of different window sizes $m$ on stylistic effects and computational costs in the sliding-window consistency smoothing scheme, we present the visual effects and corresponding inference time for various window sizes in Fig. 13. While larger window sizes yield slight visual improvements, they also result in significant inference costs. Ultimately, we select $m=2$ to balance stylistic effects and computational costs. Furthermore, we also explore how applying this strategy at different timestep intervals influences the final temporal consistency. As shown in Fig. 11, we observe that the effectiveness of the smoothing strategy varies across the early, middle, and late stages of the denoising process, with the best results achieved during the middle stage.
Figure 15: Qualitative comparisons of different inversion methods. EasyInv [38] demonstrates better transfer quality compared to DDIM-Inv [16].
5.5 Addtional Analysis
In this section, we further analyze the bottleneck of inference time, the constraints in mask propagation, and the impacts of inversion.
Bottleneck of Inference Time. In Table I and Table II, our proposed UniVST exhibits the longest inference time among all compared methods, revealing a key computational bottleneck. To better understand this limitation, we perform a detailed breakdown of the inference time distribution across the core modules of UniVST, as illustrated in Fig. 14. These modules include: DDIM Inversion, Point-Matching Mask Propagation, AdaIN-Guided Localized Video Stylization, and Sliding-Window Consistent Smoothing. Our analysis indicates that the primary source of inference overhead arises from the sliding-window consistent smoothing module. This component relies on dynamic optical flow estimation during the denoising phase, which improves temporal consistency but also adds significant computational cost. In contrast, most existing text-to-video (T2V) frameworks adopt temporal attention mechanisms, which are generally more efficient though potentially less robust in handling complex motion. To address this trade-off, a promising alternative is to replace the smoothing module with a 3D VAE decoder that directly reconstructs temporally coherent video from the denoised latent representation, thereby maintaining temporal consistency without incurring additional inference time.
Constraints in Mask Propagation. Although our point-matching mask propagation strategy enables zero-shot object localization, as demonstrated in Fig. 9, it presents several limitations: 1. Challenges in Multi-Object Scenarios. In scenes containing multiple similar objects (e.g., two individuals), the semantic similarity between instances can lead to confusion during propagation. This often results in incorrect mask associations, particularly as the propagation distance increases, causing error accumulation over time. While our current evaluation focuses on diverse yet single-object scenes—where such issues have minimal impact—extending the method to robustly handle multi-object scenarios remains an important direction for future work. 2. Limitations in Fine-Grained Mask Propagation. Although the propagation achieves an average Dice score of approximately $0.87$, it occasionally fails to capture fine object boundaries, leading to incomplete or imprecise tracking. As shown in Fig. 9, while the method effectively handles fast motion and occlusions, it struggles to retain detailed edge information. Nonetheless, our localized latent blending strategy, which performs dynamic region-wise fusion during the denoising phase, effectively mitigates this limitation by reducing reliance on high-precision masks. The mechanism can still preserve the object’s stylization quality even without guidance from high-precision masks.
Impacts of Inversion. As a training-free stylization method, UniVST relies on the initial noise derived from the original video, thereby rendering the inversion step crucial. The overall performance of UniVST is somewhat sensitive to the quality of the inverted noise, which serves as a critical foundation for subsequent stylization. Inadequate inversion can compromise both content fidelity and stylistic consistency. As illustrated in Fig. 15, we observe that naive inversion methods often fail to preserve fine-grained semantic details and adversely impact subsequent mask propagation. To mitigate the artifacts introduced by imperfect inversion, we employ EasyInv [38], a more advanced inversion technique. By leveraging EasyInv, we achieve higher-fidelity noise reconstructions, which in turn lead to improved visual quality and stable mask propagation throughout the stylization process.
Figure 16: Qualitative comparisons of our proposed UniVST with some existing commercial models. The first row presents the original video, reference image, and mask, followed by the results of Gen-1 [53] in the second row, Gen-3 in the third row, and our proposed UniVST in the final row.
Figure 17: Multi-step video style transfer: first, the background is stylized, and then the foreground is stylized based on the transformed background.
5.6 Addtional Comparison
Futhermore, we have also compared our UniVST with existing commercial models, namely Gen-1 [53] and Gen-3, as illustrated in Fig. 16. UniVST demonstrates clear superiority in two critical aspects: achieving a closer style similarity to the reference image and maintaining higher content fidelity to the original video. In addition, UniVST uniquely excels in accurately preserving the style consistency of local objects, thereby ensuring a more cohesive and visually appealing result, even in complex or dynamic scenes.
5.7 Multi-Step Video Style Transfer
Apart from the cases presented in the experimental section, UniVST showcases its versatility by enabling multi-step video style transfer, which allows for more customized transformations and the creative potential of video editing. For instance, users can start by stylizing the background of the video using a mask to isolate it from the rest of the scene. After completing this step, they can invert the mask to focus on the previously excluded foreground and apply a separate style transfer to it. This sequential process not only provides greater control but also enables creative layering of styles. As illustrated in Fig. 17, employing distinct style images for the background and foreground in these two stages produces a visually striking and dynamic effect, enhancing the artistic flexibility of the method.
6 Limitation and Future Work
While our UniVST framework shows a clear advantage in localized video style transformation, it also has limitations: (1) The accuracy of mask propagation and video content fidelity may be significantly affected by inaccuracies in the inversion process. (2) The style transfer mechanism, guided solely by prior knowledge of LDM [32], may struggle with certain style images, especially those with nuanced color distributions, leading to suboptimal performance. (3) The sliding-window smoothing technique, based on optical flow, can introduce additional inference time. Future research could explore improved stylization techniques to enhance visual effects and reduce inference overhead from optical flow prediction.
7 Conclusion
This paper presents UniVST, a unified framework for localized video style transfer without additional training, marking a significant advancement over traditional methods that require full video style transfer. Our contributions include: (1) A point-matching mask propagation strategy that eliminates the need for tracking models, enabling streamlined style transfer to specific video objects. (2) A training-free AdaIN-guided localized video stylization mechanism that operates at both the latent and attention levels, ensuring a balance between content fidelity and style richness. (3) A sliding-window consistent smoothing scheme that uses optical flow to refine noise and update the latent space, improving temporal consistency and reducing artifacts. Extensive experiments show that our UniVST outperforms existing methods in both qualitative and quantitative evaluations, effectively preserving the style of the primary object while ensuring temporal consistency. Furthermore, it can be extended to a broader range of customized models.